OpenStudy (samyra):

he needs 2 eggs to make 36 cookies,how many cookies will he need to make 90 cookies a. 2 eggs b 3 eggs c 4 eggs d 5 eggs

OpenStudy (anonymous):

d. 5 Eggs

OpenStudy (anonymous):

set up a ratio 2/36=x/90 you can reduce 2/36 to 1/18 1/18=x/90 x=90/18=5